4|-6x + 6| + -3 = 93 solve absolute value inequality

Answer by edjones(8007)

4|-6x + 6|+(-3)=93
4|-6x+6|=96
|-6x+6|=24
.
-6x+6=24
-6x=18
x=-3
.
-6x+6=-24
-6x=-30
x=5
.
Ed